15m+4-9m= -32 What Is M??
Answer by checkley71(8403)

15M+4-9M=-32 MOVE ALL Ms TO THE LEFT SIDE OF THE EQUATION & ALL NUMBERS TO THE RIGHT SIDE OF THE EQUATION THUS:
15M-9M=-32-4
6M=-36 NOW DIVIDE BOTH SIDES BY 6
6M/6=-36/6
M=-6PROOF
15*-6+4-9*-6=-32
-90+4+54=-32
-90+58=-32
-32=-32